Grade 11Mechanics

find time of flight of projectile thrown horizontally with speed 50m/s from a long inclined plane which makes an angle of 45 from horizontal.

Akshat
10 Years agoGrade 11

3 Answers

Paras Verma
10 Years ago
Horizontal distance travelled equals vertical distance travelled
Horizintal motion-  
accleration=0
velocity=50m/s(constant)     
vetical distance=x(say)
speed=distance/time
time=x/50
t2=x2/2500-------Eqn(1)
Vertical motion-
accln=g
velocity(initial)=0
s=ut+gt2/2
x=10t2/2
t2=2x/10--------Eqn(2)
from Eqn(1) and Eqn(2)
x2/2500=2x/10
x/2500=1/5
x=2500/5=500
Now t=x/50
t=500/50=10sec
 
 
Trisha Barui
8 Years ago
Dear Akshat, 
Please find the correct solution.
 
T= 2u sin theeta/g [formula for time o flight]
so t = 2*50* sin 45/10
so t= 10/ root 2
so t = 5 root 2
 please ask if any doubt.
 
Regards,
Arpit
6 Years ago
C orrect answer is root 2.